OpenStudy (anonymous):

Determine the value of constants so that the exponential form y=ae^kt passes through the points (0,3) and 93,5)

OpenStudy (anonymous):

(3,5)*

ganeshie8 (ganeshie8):

it passes through (0, 3) substitute t=0, y = 3 3 = ae^k0 => a=3

ganeshie8 (ganeshie8):

similarly use the other point and try solving for k :)

OpenStudy (anonymous):

is it a=3 and k=1/5 ln31

ganeshie8 (ganeshie8):

second point is (3,5 ) 5 = 3e^3k 5/3 = e^3k ln(5/3) = 3k k = 1/3 ln(5/3)